OpenStudy (anonymous):

What is the sum of the first 28 terms of this arithmetic sequence? 69, 75, 81, 87, 93, 99, … A. 2,436 B. 4,200 C. 4,320 D. 4,368

OpenStudy (chrisdbest):

B

OpenStudy (chrisdbest):

I actually wrote them all down, sorry.

OpenStudy (freckles):

\[S_n=\frac{n(a_1+a_n)}{2}\] Sn is the sum of the first n terms a1 is the first term an is the nth term

OpenStudy (freckles):

you can find the 28 term by using: \[a_n=a_1+d(n-1) \\ a_{28}=69+d(28-1)\] I will let you say what d is

OpenStudy (anonymous):

THANKS